# Sparse Subnetwork Enhancement for Underrepresented Languages in Large Language Models

### v1

@@ Abstract
Large language models exhibit uneven performance across languages, with substantial gaps between high- and low-resource languages. We present a framework for enhancing monolingual capabilities of LLMs in underrepresented languages while preserving their general-purpose performance through targeted fine-tuning of language-specific subnetworks. Our approach identifies language-specific neurons using Language Activation Probability Entropy and fine-tunes only the weights associated with these neurons, a dedicated subnetwork, on target-language data. Experiments on Llama-3.1-8B and Mistral-Nemo-12B across 12 mid- and low-resource languages demonstrate that our method consistently outperforms full fine-tuning, FFN-only fine-tuning, LoRA adaptation, and random subset fine-tuning baselines while efficiently updating only up to 1% of model parameters. Beyond performance improvements, we observe enhanced favorable training dynamics, cross-lingual representational alignment, and systematic weight update changes. To facilitate future research, we release language-specific neuron identifications for over 100 languages as well as our adaptation pipeline, offering a cost-effective pathway for adapting state-of-the-art models to underrepresented languages.
Sparse Subnetwork Enhancement for Underrepresented Languages
in Large Language Models
Daniil Gurgurov1,2  Josef van Genabith1,2  Simon Ostermann1,2,3 1Saarland University 2German Research Center for Artificial Intelligence (DFKI) 3Centre for European Research in Trusted AI (CERTAIN) daniil.gurgurov@dfki.de
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse tasks and languages, yet their performance remains unevenly distributed, with substantial gaps between high-resource and low-resource languages Joshi et al. (2020); Huang et al. (2023). While scaling training data and model parameters has improved multilingual performance, this approach is still often impractical and ineffective for resource-constrained scenarios involving hundreds of languages Robinson et al. (2023).
@@
Figure 1: Sparse subnetwork adaptation method: attention blocks remain frozen (grey) while selected language-specific FFN neurons and their associated weights are fine-tuned (red). This approach enables parameter-efficient adaptation for underrepresented languages while preserving the model’s general knowledge.
@@
Recent advances in mechanistic interpretability have revealed that language models develop specialized internal structures for different languages, with distinct neurons and attention patterns activated for specific linguistic contexts Tang et al. (2024); Zhao et al. (2024); Kojima et al. (2024). This observation suggests a promising alternative to full model retraining: identifying and selectively enhancing language-specific components while preserving general-purpose capabilities.
@@
Current approaches to multilingual adaptation face a fundamental trade-off between target-language performance and retention of general capabilities. Full fine-tuning often leads to catastrophic forgetting French (1999), while parameter-efficient methods like adapters Houlsby et al. (2019a), LoRA Hu et al. (2021), and soft prompts Vykopal et al. (2025) may not fully exploit the model’s existing multilingual structure. Existing neuron-level approaches have shown promise but have primarily focused on task-specific adaptation rather than systematic language enhancement, and they have often fallen short in preserving existing knowledge Zhu et al. (2024); Xu et al. (2024); Mondal et al. (2025).
@@
In this work, we present a framework for enhancing monolingual capabilities of LLMs through targeted fine-tuning of language-specific subnetworks within feed-forward network (FFN) components. Our approach leverages Language Activation Probability Entropy (LAPE) Tang et al. (2024) to identify neurons that exhibit strong language-specific activation patterns within Llama-3.1-8B and Mistral-Nemo-12B, then fine-tunes only the weights associated with these neurons using target-language data. This sparse fine-tuning updates up to 1% of model parameters while achieving substantial improvements in target-language performance and preserving general capabilities.
@@
  • We introduce a systematic framework for efficient and effective language subnetwork identification and enhancement that preserves general-purpose performance while significantly improving target-language capabilities.
  • We demonstrate the effectiveness of our approach across 12 mid- and low-resource languages, showing consistent improvements over full fine-tuning, FFN-only fine-tuning, LoRA fine-tuning, and random subset fine-tuning baselines.
  • We show that target-language performance can be improved by increasing the amount of training data where possible, laying the foundation for optimal adaptation strategies across different languages.
  • Through activation analysis, we show that targeted subnetwork fine-tuning improves cross-lingual alignment, while weight-change analysis reveals that its effects are concentrated in the later down-projection matrices of FFN layers, highlighting their central role in adapting to new languages.
  • We release language-specific neuron identifications for over 100 languages, providing a foundation for future research in low-resource language adaptation. [^1] https://github.com/d-gurgurov/Language-Subnetwork-Enhancement-LLMs.
This work represents a step toward democratizing high-quality language model capabilities across diverse languages. By providing both a practical framework and comprehensive resources, we aim to lower the barrier for adapting state-of-the-art models to underrepresented languages.
@@ 2.1 Full Fine-Tuning Approaches
Full fine-tuning is the most direct strategy for language adaptation. Post-training on related languages has been used to mitigate overfitting Neubig and Hu (2018), while domain-adaptive fine-tuning improved contextualized models on specialized corpora Han and Eisenstein (2019). Language-specific fine-tuning on monolingual corpora Gururangan et al. (2020); Chau et al. (2020), transliteration-based adaptation Muller et al. (2021), and even small corpora such as Bible translations Ebrahimi and Kann (2021) have boosted performance on tagging, parsing, and NER. However, full fine-tuning is costly, data-hungry, and prone to catastrophic forgetting French (1999), motivating parameter-efficient alternatives.
2.2 Adapter-Based Approaches
Adapters Houlsby et al. (2019b) introduce small trainable modules into frozen backbones, reducing computational cost Strubell et al. (2019) and data requirements Faisal and Anastasopoulos (2022) while avoiding forgetting French (1999). MAD-X Pfeiffer et al. (2020), UDapter Üstün et al. (2020), and MAD-G Ansell et al. (2021) extend adapters with task/language modularity and typological signals, while other work explores hierarchical designs Faisal and Anastasopoulos (2022), language balancing Lee et al. (2022); Parović et al. (2022), and inference-aware training Parović et al. (2023). Recent studies show adapter tuning can outperform continued pre-training for adaptation Yong et al. (2023); Gurgurov et al. (2025b).
@@ 2.3 Neuron-Level Approaches
Recent work has explored fine-grained adaptation by targeting individual neurons or attention heads. Zhu et al. (2024) identify language-aware neurons in FFNs using Taylor scores Xie et al. (2021) and fine-tune only the relevant subsets using parallel data for machine translation, outperforming both adapters and full fine-tuning in high-resource settings. Similarly, Zhao et al. (2024) adapt $\sim$0.1% of neurons in FFNs and attention heads identified via neuron ablations, observing initial gains for a subset of languages but rapid saturation with additional data. Xu et al. (2024) propose NeFT, which compares pre- and post-fine-tuning representations to locate task-sensitive neurons, and show that fine-tuning just 5–10% of them consistently surpasses LoRA Hu et al. (2021) on translation and summarization. Mondal et al. (2025) apply LAPE Tang et al. (2024) for neuron identification and manipulation, but report limited improvements in cross-lingual transfer. Their setup, however, differs from ours in that they restrict analysis to only three relatively high-resource languages and adapt LoRA modules on top of FFNs rather than directly fine-tuning neurons inside the model, which makes their findings not directly comparable to ours. Complementary work shows that only a few attention heads encode translation mechanisms, and selectively adapting these heads matches full fine-tuning while preserving general abilities Zhang et al. (2025).
2.4 Towards Targeted Adaptation
Neuron- and head-level methods demonstrate that precise adaptation can rival or exceed large-scale fine-tuning, yet prior work has emphasized task-specific gains rather than systematic language adaptation. In particular, it remains unclear to what extent language-specific information can be injected into carefully chosen subnetworks for underrepresented languages without harming general performance. Our work addresses this gap by leveraging LAPE, an information-theoretic method that has recently shown strong potential for precise language neuron identification Tang et al. (2024) and manipulation Gurgurov et al. (2025a), to identify language-specific subnetworks and fine-tune their associated weights on target-language data.
@@ 3 Methodology
Our approach consists of two components: (i) identifying language-sensitive neurons, and (ii) selectively fine-tuning the corresponding subnetwork (see Figure 1). Together, these steps allow us to adapt LLMs to underrepresented languages while minimizing changes to unrelated parameters.
@@ 3.1 Neuron Identification
We leverage Language Activation Probability Entropy (LAPE) Tang et al. (2024) to locate neurons that exhibit language-specific behavior. We focus on neurons within FFN components, since prior work has shown that they frequently encode rich factual knowledge Geva et al. (2021); Dai et al. (2021); Wang et al. (2022) and multilingual information Zhao et al. (2024).
For each neuron $j$ in layer $i$, we compute the probability of activation under inputs from language $k$:
$p^{k}_{i,j}=\mathbb{E}_{x\sim\mathcal{D}_{k}}\Big[\mathbf{1}\!\left(\phi(h_{i}(x)W^{i}_{1})_{j}>0\right)\Big],$
where $h_{i}(x)$ is the hidden state entering layer $i$, $W^{i}_{1}$ is the input projection of the feed-forward network, $\phi$ is the activation function, and $\mathbf{1}$ is the indicator function.
The vector of probabilities across all languages $k=1,\dots,L$ is normalized:
$\tilde{p}_{i,j}^{k}=\frac{p^{k}_{i,j}}{\sum_{\ell=1}^{L}p^{\ell}_{i,j}}.$
The LAPE score is then given by Shannon entropy Shannon and Weaver (1998):
$\text{LAPE}_{i,j}=-\sum_{k=1}^{L}\tilde{p}_{i,j}^{k}\log\tilde{p}_{i,j}^{k}.$
We select neurons with the lowest $K\%$ of LAPE scores, indicating truly language-specific behavior due to their low-entropy activations, subject to two conditions:
1. $\max_{k}p^{k}_{i,j}\geq\tau_{\text{activity}}$: a neuron must reach at least $\tau_{\text{activity}}$ in some language; otherwise, uniformly weak neurons are discarded.
2. $\exists k:p^{k}_{i,j}\geq\tau_{\text{selectivity}}$: a neuron is kept only if its activation for at least one language exceeds a strong threshold $\tau_{\text{selectivity}}$, ensuring clear language preference.
The resulting subset forms a language-sensitive subnetwork $\mathcal{S}^{k}$.
@@ 3.2 Subnetwork Fine-Tuning
Given the identified subnetwork $\mathcal{S}^{k}$, we restrict parameter updates to the weights associated with these neurons. Let $\theta$ denote the full set of model parameters, and $\theta_{\mathcal{S}^{k}}\subset\theta$ the parameters belonging to the selected neurons.
Fine-tuning is then performed by minimizing the loss $\mathcal{L}$ on target-language data $\mathcal{D}_{\text{target}}$:
$\min_{\theta_{\mathcal{S}^{k}}}\;\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{target}}}\;\mathcal{L}\big(f(x;\theta),y\big),$
while keeping all parameters $\theta\setminus\theta_{\mathcal{S}^{k}}$ fixed.
This optimization aims at adapting only the language-relevant subnetwork while preserving the general-purpose behavior of the model.

### v2

@@ Abstract
Large language models (LLMs) exhibit substantial performance disparities across languages, particularly between high- and low-resource settings. We propose a framework for improving performance in underrepresented languages while preserving general-purpose capabilities via targeted fine-tuning of sparse, language-associated subnetworks. Our approach identifies language-relevant neurons using Language Activation Probability Entropy (LAPE), an information-theoretic metric that reliably captures language-specific activation patterns, and fine-tunes only the corresponding weights. Experiments on Llama-3.1-8B, Mistral-Nemo-12B, and Aya-Expanse-8B across 12 mid- and low-resource languages show that our method consistently outperforms full fine-tuning, FFN-only fine-tuning, LoRA, IA3, and random-subset baselines while updating only 0.2–1% of model parameters. We further show that sparse, neuron-targeted fine-tuning can inject new language capabilities without catastrophic forgetting, with potential applicability to other model capabilities. Mechanistic analyses of weight updates and internal representations reveal asymmetric roles of FFN projections in language adaptation and improved cross-lingual alignment. Finally, we release language neuron sets for over 100 languages together with our adaptation pipeline, enabling a cost-effective path for extending LLMs to underrepresented languages.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse tasks and languages, yet their performance remains unevenly distributed, with substantial gaps between high-resource and low-resource languages (Joshi et al., 2020; Huang et al., 2023). While scaling training data and model parameters has improved multilingual performance, this approach is often impractical for resource-constrained scenarios involving hundreds of languages (Robinson et al., 2023; Lai et al., 2024).
Multi-HeadAttentionFeed-Forward NetworkTrainedFrozen
@@
Figure 1: Sparse subnetwork adaptation. Attention blocks remain frozen while LAPE-identified language-specific neurons and their weights within FFNs are fine-tuned.
@@
Recent advances in mechanistic interpretability have revealed that language models develop specialized internal structures for different languages, with distinct neurons and attention patterns activated for specific linguistic contexts (Tang et al., 2024; Zhao et al., 2024; Liu et al., 2024; Kojima et al., 2024). This observation suggests a promising alternative to full model retraining: identifying and selectively enhancing language-specific components while preserving general-purpose capabilities.
@@
Current approaches to multilingual adaptation face a fundamental trade-off between target-language performance and retention of general capabilities. Full fine-tuning often leads to catastrophic forgetting (French, 1999; Kirkpatrick et al., 2017), while parameter-efficient methods (PEFT) (Houlsby et al., 2019; Hu et al., 2022; Vykopal et al., 2025) have emerged as practical alternatives; however, they operate without explicit consideration of how knowledge is organized within the model, which may lead to the indiscriminate modification of parameters encoding different functions. Neuron-level approaches have shown promise in this regard but often underperform PEFT methods. Moreover, they remain largely unexplored for systematic language enhancement: prior work has focused on task-specific adaptation, such as machine translation (Zhu et al., 2024; Xu et al., 2025), or employed neuron identification primarily for interpretability rather than targeted fine-tuning (Tang et al., 2024; Kojima et al., 2024; Wang et al., 2025). A key question remains open: can we inject new language capabilities into LLMs by directly fine-tuning only the sparse subset of neurons that encode language-specific processing?
@@
In this work, we provide evidence that this is indeed feasible. We present a framework for enhancing monolingual capabilities of LLMs through targeted fine-tuning of language-specific subnetworks within feed-forward network (FFN) components. We focus on FFNs, which prior work has shown to be the primary site of linguistic knowledge in transformers (Geva et al., 2021; Wang et al., 2022). Our approach uses Language Activation Probability Entropy (LAPE) (Tang et al., 2024), an information-theoretic metric that quantifies how selectively neurons activate for specific languages, to identify neurons exhibiting strong language-specific patterns within Llama-3.1-8B (Grattafiori et al., 2024), Mistral-Nemo-12B (Mistral AI Team, 2024), and Aya-Expanse-8B (Dang et al., 2024). We then fine-tune only the weights associated with these neurons using target-language data, updating just 0.2–1% of model parameters.
Our central finding is that sparse, targeted fine-tuning of language-specific neurons can systematically inject new language capabilities without catastrophic forgetting while outperforming common PEFT methods, a result that, to our knowledge, has not been demonstrated at this scale. Beyond performance gains, we provide mechanistic insights into how language adaptation occurs: weight-change analysis reveals asymmetric roles of FFN projection matrices, with down-projections in later layers showing substantially larger updates, suggesting they serve as the primary site of linguistic re-wiring during adaptation.
@@
  • We introduce a systematic framework for enhancing underrepresented languages through sparse subnetwork fine-tuning, demonstrating that language capabilities can be injected by updating less than 1% of parameters.
  • We provide empirical validation across 12 mid- and low-resource languages on three models, showing (i) consistent absolute improvements over full fine-tuning, FFN-only fine-tuning, LoRA, IA3 (Liu et al., 2022), and random subset baselines and (ii) full preservation of general capabilities, which is otherwise only exhibited by IA3.
  • We offer mechanistic insights into language adaptation: cross-lingual alignment improves systematically, and weight changes concentrate in later-layer down-projections, revealing their central role in integrating new language-specific knowledge.
  • We release language-specific neuron identifications for over 100 languages alongside our adaptation pipeline, providing practical resources for the community. [^1] Code, neuron indices, and adapted models are available on GitHub and Hugging Face.
This work represents a step toward democratizing high-quality language model capabilities. By demonstrating that principled subnetwork identification enables effective adaptation with minimal computational resources, we provide a practical framework and a foundation for unlocking future research on LLMs for underrepresented languages.
@@ 2.1 Full Fine-Tuning Approaches
Full fine-tuning is the most direct strategy for language adaptation. Post-training on related languages has been used to mitigate overfitting (Neubig & Hu, 2018), while domain-adaptive fine-tuning improved contextualized models on specialized corpora (Han & Eisenstein, 2019). Language-specific fine-tuning on monolingual corpora (Gururangan et al., 2020; Chau et al., 2020), transliteration-based adaptation (Muller et al., 2021), and even small corpora such as Bible translations (Ebrahimi & Kann, 2021) have boosted performance on tagging, parsing, and NER. However, full fine-tuning is costly, data-hungry, and prone to catastrophic forgetting (French, 1999), motivating parameter-efficient alternatives.
2.2 Parameter-Efficient Approaches
Adapters (Houlsby et al., 2019) introduce small trainable modules into frozen backbones, substantially reducing computational cost (Strubell et al., 2019) and data requirements while mitigating catastrophic forgetting (French, 1999). Prominent multilingual extensions such as MAD-X (Pfeiffer et al., 2020), UDapter (Üstün et al., 2020), and MAD-G (Ansell et al., 2021) incorporate language- and task-specific modularity, typological signals, or hierarchical structure, while other work explores language balancing (Lee et al., 2022a; Parović et al., 2022) and inference-aware training (Parovic et al., 2023). Recent studies further show that adapter tuning can outperform continued pre-training for language adaptation (Yong et al., 2023; Gurgurov et al., 2025b). Beyond adapters, several parameter-efficient alternatives have been proposed. IA3 (Liu et al., 2022) scales activations using learned vectors, Compacter (Karimi Mahabadi et al., 2021) employs low-rank hypercomplex adapters, and AdaLoRA (Zhang et al., 2023) dynamically allocates rank budgets during training. While effective for task adaptation, these approaches introduce auxiliary parameters and do not explicitly leverage or modify the model’s internal multilingual structure.
@@ 2.3 Neuron-Level Approaches
Recent work has explored fine-grained adaptation by targeting individual model components. Early approaches used gradient-based criteria, such as magnitude pruning (Han et al., 2015), movement pruning (Sanh et al., 2020), Taylor expansion scores (Molchanov et al., 2017), to identify important parameters; though these were developed for older CNN and encoder-only architectures. For modern LLMs, SLAM (Fan et al., 2025) selects layers with high multilingual activation variance and fine-tunes their FFN sublayers, but operates at layer granularity. Zhu et al. (2024) adapt Taylor scores to identify language-aware components for machine translation, but require gradient computation on parallel data, limiting applicability to low-resource settings. Zhao et al. (2024) use ablation studies to identify neurons and fine-tune $\sim$0.1% of parameters for four high-resource languages, but report rapid saturation with additional data due to overfitting to the training source. Xu et al. (2025) propose NeFT, identifying task-sensitive neurons by comparing pre- and post-fine-tuning representations, though this requires a full fine-tuning pass before neuron selection. Mondal et al. (2025) apply LAPE for neuron identification but observe limited gains; notably, their study focuses on three high-resource languages and adapts LoRA modules atop FFNs rather than directly fine-tuning neuron weights. previous work has not succeeded in using these structures reliably
2.4 Positioning of Our Work
Despite growing evidence that language-specific structures exist inside LLMs, previous work has not succeeded in using these structures reliably to add new language capabilities at scale without degrading existing ones. Existing neuron identification methods rely on task-specific signals: Taylor scores require gradient computation on parallel data (Zhu et al., 2024), NeFT requires a fine-tuning pass to identify task-sensitive neurons (Xu et al., 2025), and ablation-based methods (Zhao et al., 2024) are computationally expensive at scale. In contrast, LAPE (Tang et al., 2024) offers a task-agnostic, information-theoretic measure of language selectivity derived directly from activation statistics and requires only monolingual text. Prior work has shown that LAPE enables robust neuron identification and target manipulation across languages (Tang et al., 2024; Gurgurov et al., 2025a). Building on this foundation, we establish sparse subnetwork fine-tuning as a viable mechanism for capability-level language adaptation, rather than task-specific adjustment, while explicitly maintaining general model competence. We demonstrate that this approach holds across models, languages, and benchmarks.
@@ 3 Methodology
Our approach consists of two parts: (i) identifying language-sensitive neurons, and (ii) selectively fine-tuning the corresponding subnetwork (Figure 1). This enables adaptation to underrepresented languages while minimizing updates to unrelated parameters.
@@ 3.1 Neuron Identification
We leverage Language Activation Probability Entropy (LAPE) (Tang et al., 2024) to identify neurons exhibiting language-specific behavior. As discussed in Section 2, LAPE is task-agnostic and requires only monolingual text, making it well-suited to low-resource settings. We focus on neurons within FFN components, as prior work has established that feed-forward layers serve as the primary focus of factual and linguistic knowledge storage (Geva et al., 2021; Dai et al., 2022; Meng et al., 2022), whereas attention heads primarily mediate information routing and contextual dependencies (Voita et al., 2019; Clark et al., 2019).
We compute LAPE over a reference set of $K=12$ typologically diverse underrepresented languages (Joshi et al., 2020): Maltese, Afrikaans, Icelandic, Welsh, Macedonian, Latvian, Lithuanian, Slovenian, Slovak, Estonian, Georgian, and Nepali.
For each neuron $i$ in layer $\ell$, we estimate the probability of activation under inputs from language $k$:
$p^{k}_{\ell,i}=\mathbb{E}_{x\sim\mathcal{D}_{k}}\Big[\mathbf{1}\!\left(\sigma(W_{\text{gate}}^{(\ell)}h^{(\ell)}(x))_{i}>0\right)\Big],$ (1)
where $\mathcal{D}_{k}$ is a monolingual corpus for language $k$, $h^{(\ell)}(x)\in\mathbb{R}^{d}$ is the hidden state input to layer $\ell$, $W_{\text{gate}}^{(\ell)}\in\mathbb{R}^{d_{\text{ff}}\times d}$ is the gate projection matrix, $\sigma(\cdot)$ is the SiLU activation function, and $\mathbf{1}(\cdot)$ is the indicator function.
The probability vector across all languages $k\in\{1,\dots,K\}$ is normalized to obtain a distribution:
$\tilde{p}_{\ell,i}^{k}=\frac{p^{k}_{\ell,i}}{\sum_{k^{\prime}=1}^{K}p^{k^{\prime}}_{\ell,i}}.$ (2)
The LAPE score is then computed as the Shannon entropy (Shannon & Weaver, 1998) of this distribution:
$\text{LAPE}_{\ell,i}=-\sum_{k=1}^{K}\tilde{p}_{\ell,i}^{k}\log\tilde{p}_{\ell,i}^{k}.$ (3)
Low LAPE scores indicate language-specific neurons, as their activations concentrate on a small subset of languages. We select neurons in the lowest $\rho$-th percentile of LAPE scores ($\rho=5\%$), subject to an activity threshold $\tau_{\text{act|}|}$, defined as the 95th percentile of activation probabilities across all neurons and languages. This threshold is applied in two ways: (1) globally, we require $\max_{k}p^{k}_{\ell,i}\geq\tau_{\text{act}}$ to discard uniformly inactive neurons that achieve low entropy trivially; (2) per-language, we assign a neuron to the subnetwork $\mathcal{S}^{k}$ only if $p^{k}_{\ell,i}\geq\tau_{\text{act|}|}$.
Since assignment is based on exceeding the activation threshold rather than exclusive assignment to the maximally-activating language, neurons with high activation across multiple languages may belong to multiple subnetworks. This reflects the linguistic reality that some neurons encode features shared across related languages.
@@ 3.2 Subnetwork Fine-Tuning
Given the subnetwork $\mathcal{S}^{k}$ for language $k$, we restrict parameter updates to the weights associated with the selected neurons. We consider transformer architectures with GLU-based feed-forward blocks (Shazeer, 2020), which contain three weight matrices per layer $\ell$:
  • Gate projection: $W_{\text{gate}}^{(\ell)}\in\mathbb{R}^{d_{\text{ff}}\times d}$
  • Up projection: $W_{\text{up}}^{(\ell)}\in\mathbb{R}^{d_{\text{ff}}\times d}$
  • Down projection: $W_{\text{down}}^{(\ell)}\in\mathbb{R}^{d\times d_{\text{ff}}}$
where $d$ is the hidden dimension and $d_{\text{ff}}$ is the intermediate dimension. The FFN computation for input $\mathbf{h}\in\mathbb{R}^{d}$ is given by:
$\text{FFN}(\mathbf{h})=W_{\text{down}}\cdot\Big(\sigma(W_{\text{gate}}\mathbf{h})\odot(W_{\text{up}}\mathbf{h})\Big),$ (4)
where $\sigma(\cdot)$ denotes the SiLU activation function and $\odot$ represents element-wise multiplication.
For each neuron $i$ in the subnetwork $\mathcal{S}^{k,(\ell)}$ at layer $\ell$, we define the associated trainable parameters as:
$\theta_{i}^{(\ell)}=\left\{W_{\text{gate}}^{(\ell)}[i,:],\;W_{\text{up}}^{(\ell)}[i,:],\;W_{\text{down}}^{(\ell)}[:,i]\right\},$ (5)
where $W[i,:]$ denotes the $i$-th row and $W[:,i]$ denotes the $i$-th column. The complete set of trainable parameters for language $k$ is then:
$\theta_{\mathcal{S}^{k}}=\bigcup_{\ell=1}^{L}\bigcup_{i\in\mathcal{S}^{k,(\ell)}}\theta_{i}^{(\ell)}.$ (6)
Fine-tuning is performed by minimizing the language modeling loss on target-language data $\mathcal{D}_{\text{target}}^{k}$:
$\min_{\theta_{\mathcal{S}^{k}}}\;\mathbb{E}_{x\sim\mathcal{D}_{\text{target}}^{k}}\left[-\sum_{t=1}^{|x|}\log P(x_{t}\mid x_{<t};\theta)\right],$ (7)
while keeping all remaining parameters $\theta\setminus\theta_{\mathcal{S}^{k}}$ frozen. This selective optimization adapts only the language-relevant subnetwork while preserving the model’s general capabilities.